# Limit Theorems and Quantitative Statistical Stability for the Equilibrium States of Piecewise Partially Hyperbolic Maps

### v1

@@ 7.1. An example of $\mathcal{S}$ and $\overline{\mathcal{P}}$
If the fiber $K$ can be decomposed as a finitely union $K=K_{1}\cup\cdots\cup K_{n}$ of pairwise disjoint compact sets $K_{1},\cdots,K_{n}$ then the condition $G(x,y_{0})=y_{0}$ for all $x\in M$ in the above definition can be replaced by $G_{i}(x,y_{i})=y_{i}$ for all $x\in M$ and some $y_{i}\in K_{i}$, $i=1,\cdots,n$. In fact, since $M\times K$ is a product space and $K=K_{1}\cup\cdots\cup K_{n}$ we may define $n$ fiber dynamics $G_{i}:M\times K_{i}\to K_{i}$ by $G_{i}(x,y)=G(x,y)$ when $y\in K_{i}$, for each $i=1,\cdots,n$.
In what follows $\operatorname{H}_{\zeta}(\Sigma)$ denotes the space of real $\zeta$-Hölders functions defined on $\Sigma:=M\times K$.
Given $\Phi\in\operatorname{H}_{\zeta}(\Sigma)$ and a fixed element $y_{0}\in K$ (defined in Equation (53)), we denote by $\Phi_{y_{0}}$ the function $\Phi_{y_{0}}:M\to\mathbb{R}$ defined by $\Phi_{y_{0}}(x):=\Phi(x,y_{0})$.
Definition 7.2.
Define the set of potentials $\overline{\mathcal{P}}$ by
$\overline{\mathcal{P}}:=\{\Phi\in\operatorname{H}_{\zeta}(\Sigma)|\Phi_{y_{0}}\in\mathcal{PH}\}.$ (54)
We say that two potentials $\bar{\Phi},\Phi:M\times K\to\mathbb{R}$ are homologous if there exists a continuous function $u:M\times K\to\mathbb{R}$ such that
$\bar{\Phi}=\Phi-u+u\circ F.$
The following proposition is taken from [19] (more precisely, it is part of Proposition 5.1 therein). Throughout, we assume that $F\in\mathcal{S}$ and that $F$ is continuous.
@@
Let $\Phi:M\times K\to\mathbb{R}$ be a Hölder continuous potential. There exists a Hölder continuous potential $\bar{\Phi}:M\times K\to\mathbb{R}$ not depending on the stable direction such that:
@@
Equation (8) of [19] ensures that, if $F$ is continuous, $F\in\mathcal{S}$ and $\Phi:M\times K\to\mathbb{R}$ is Hölder continuous, then $\bar{\Phi}(x,y)=\Phi(x,y_{0})$ for all $x\in M$ and $y_{0}$ is from Equation (53). Thus, $\bar{\Phi}=\Phi_{y_{0}}\circ\pi_{1}$, where $\Phi_{y_{0}}:M\longrightarrow\mathbb{R}$ is defined by $\Phi_{y_{0}}(x):=\Phi(x,y_{0})$ for all $x\in M$. Therefore, if $\Phi\in\overline{\mathcal{P}}$, then $\bar{\Phi}$ given by Proposition 7.3 is of the type
$\bar{\Phi}=\Phi_{y_{0}}\circ\pi_{1},\ \text{where}\ \Phi_{y_{0}}\in\mathcal{P}.$ (55)
The next example was introduced in [15], where it was shown that the non-wandering set of $F$ is partially hyperbolic for certain fixed parameters satisfying the conditions stated below. This class of maps was subsequently studied in [16], [18], [17], and [19]. In [18], the inverse map $F^{-1}$ was considered and shown to admit a skew product structure, where the base dynamics is strongly topologically mixing and non-uniformly expanding, while the fiber dynamics is uniformly contracting. In [19], the author established the existence of equilibrium states and proved stability results for this class of systems. It is important to note that $F$ is a diffeomorphism belonging to the class $\mathcal{S}$ (see Equation (53)).
Example 7.1.
(Partially hyperbolic horseshoes) Consider the cube $R=[0,1]\times[0,1]\times[0,1]\subset\mathbb{R}^{3}$ and the paralellepipeds
Consider a map defined for $(x,y,z)\in R_{0}$ as
$F_{0}(x,y,z)=(\rho x,f(y),\beta z),$
where $0<\rho<{1/3}$, $\beta>6$ and
$f(y)=\frac{1}{1-\left(1-\frac{1}{y}\right)e^{-1}}.$
Consider also a map defined for $(x,y,z)\in R_{1}$ as
where $0<\sigma<{1/3}$ and $3<\beta_{1}<4$. Define the horseshoe map $F$ on $R$ as
$F|_{R_{0}}=F_{0},\quad F|_{R_{1}}=F_{1},$
Remark 7.5.
Other classes $\mathcal{S}$ and $\overline{\mathcal{P}}$ can also be constructed. For instance, see Proposition 7.1 in [20], where the authors obtain a result analogous to Proposition 7.3 for homeomorphisms.

### v2

@@ 7.1. An example of $\mathcal{S}$ and $\overline{\mathscr{P}}_{\Sigma}$
If the fibre $K$ can be decomposed as a finitely union $K=K_{1}\cup\cdots\cup K_{n}$ of pairwise disjoint compact sets $K_{1},\cdots,K_{n}$ then the condition $G(x,y_{0})=y_{0}$ for all $x\in M$ in the above definition can be replaced by $G_{i}(x,y_{i})=y_{i}$ for all $x\in M$ and some $y_{i}\in K_{i}$, $i=1,\cdots,n$. In fact, since $M\times K$ is a product space and $K=K_{1}\cup\cdots\cup K_{n}$ we may define $n$ fibre dynamics $G_{i}:M\times K_{i}\to K_{i}$ by $G_{i}(x,y)=G(x,y)$ when $y\in K_{i}$, for each $i=1,\cdots,n$.
We say that two potentials $\bar{\Phi},\Phi:M\times K\to\mathbb{R}$ are co-homologous if there exists a continuous function $u:M\times K\to\mathbb{R}$ such that
$\Phi=\bar{\Phi}-u+u\circ F.$
@@
Let $\bar{\Phi}:M\times K\to\mathbb{R}$ be a Hölder continuous potential. Then there exists a Hölder continuous potential $\Phi:M\times K\to\mathbb{R}$, independent of the stable direction, such that:
@@
By Equation (8) in [19], if $F$ is continuous, satisfies (53) for some $y_{0}\in K$, and $\bar{\Phi}:M\times K\to\mathbb{R}$ is Hölder continuous, then the fibre-constant potential $\Phi$ defined by
$\Phi(x,y):=\bar{\Phi}(x,y_{0}),\qquad\forall(x,y)\in M\times K,$
is co-homologous to $\bar{\Phi}$. This is precisely the potential whose existence is guaranteed by Proposition 7.2. In particular, the pairs $(F,\Phi)$ and $(F,\bar{\Phi})$ have the same equilibrium states.
Moreover, $\Phi$ can be written as
$\Phi=\bar{\varphi}_{y_{0}}\circ\pi_{1},$
where $\bar{\varphi}_{y_{0}}:M\to\mathbb{R}$ is defined by
$\bar{\varphi}_{y_{0}}(x):=\bar{\Phi}(x,y_{0}),\qquad\forall x\in M.$
Therefore, provided that $\bar{\varphi}_{y_{0}}\in\mathscr{P}_{M}$, the class $\mathcal{S}$ defined above provides a family of systems to which all the results of this article can be applied, specifically for the set of potentials $\overline{\mathscr{P}}_{\Sigma}$ that are not constant along fibers.
@@
Other classes $\mathcal{S}$ and $\overline{\mathscr{P}}_{\Sigma}$ can also be constructed. For instance, see Proposition 7.1 in [20], where the authors obtain a result analogous to Proposition 7.2 for homeomorphisms.